Construct the probability distribution for the random variables describe each following situations. Draw the corresponding histogram for each probability distribution. Label it properly

1. Let T be a random variable giving the number of tails in three tosses of a coin. List the elements

of the sample space S for the three tosses of the coin. And assign a value to each sample point.


Valueof the Random Variable T

(Number of Tails)


Probability P(T)


1
Expert's answer
2021-04-15T07:21:21-0400

Sample space S={HHH,HHT,HTH,HTT,THH,THT,TTH,TTT}


There are total 8 sample space


The value of T can be 0,1,2,3


"P(T=0)=\dfrac{1}{8}"


"P(T=1)=\dfrac{3}{8}"


"P(T=2)=\dfrac{3}{8}"


"P(T=3)=\dfrac{1}{8}"


Probability Distribution is-





and Histogram is-